a) If X1 and X2 are independent random variables and X1 tollows the Nor nalLA σ1 X, +X2 follow? di tri t on and X to ows the Nonna μα 2 distribution, ne ha distribution do b) IfX1 , X2 . X, , arendependent random variables and each Xk follows the NormalA 에 ds rbutio. then what distribution does follow? , n L.6) Generating functions for sums of independent random variables a) If X and X are independent random variables, then how is the generating function of Xi X2 related to the generating function of X1 and the generating function of X? b) If X1 , X2 X3 .. X are independent random variables, then how do you get the generating function of Xi +X2+X3+...X from the generating functions of XiX^ .Xs ....X-i .and X?
Answer #1

a)

We know that the linear combination of Normal random variables follow Normal distribution.

E[X1 + X2] = E[X1] + E[X2] = mu_1 + mu_2

Var[X1 + X2] = Var[X1] + Var[X2] = 2

Standard deviation of X1 + X2 is 2 2

Thus,

X1 + X2 follows the Normal[mu_1 + mu_2, 2 2 ]

b)

We know that the linear combination of Normal random variables follow Normal distribution.

に! k=1 に!

に! k=1 に! (Since all Xk are independent random variables)

Standard deviation of sum_{k=1}^{n}X_k is,

sqrt{sum_{k=1}^{n} sigma^2_k ight}

Thus, sum_{k=1}^{n}X_k follows the

7L 7L

a)

Moment generating function of X1 + X2 is product of moment generating function of X1 and X2.

M_{X_1+X_2} (t)= M_{X_1} (t).M_{X_2} (t)

b)

Moment generating function of X1 + X2 + ... + Xn is product of moment generating function of X1, X2, ..., Xn-1 , Xn.

M_{X_1+X_2+ ..+ X_n} (t)= M_{X_1} (t).M_{X_2} (t)...M_{X_{n-1}} (t).M_{X_n} (t)